\begin{thebibliography}{Fekete2020}

\bibitem[AOC12]{AOC}
Y.~Ameur and J.~Ortega-Cerd\`a
\newblock Beurling-Landau densities of weighted Fekete sets and
correlation kernel estimates.
\newblock {\em J. Funct. Anal.}, 263: 1825--1861, 2012.


\bibitem[Bay16]{B6}
T.~Bayraktar.
\newblock Equidistribution of zeros of random holomorphic sections.
\newblock {\em Indiana Univ. Math. J.}, 65(5):1759--1793, 2016.

\bibitem[Bay17]{B4}
T.~Bayraktar.
\newblock Asymptotic normality of linear statistics of zeros of random
  polynomials.
\newblock {\em Proc. Amer. Math. Soc.}, 145(7):2917--2929, 2017.

\bibitem[Bay20]{B8}
T.~Bayraktar.
\newblock Mass equidistribution for random polynomials.
\newblock {\em Potential Anal.}, DOI:10.1007/s11118-019-09811-w.

\bibitem[BCHM18]{BCHM}
T.~Bayraktar, D.~Coman, H.~Herrmann, and G.~Marinescu.
\newblock A survey on zeros of random holomorphic sections.
\newblock {\em Dolomites Res. Notes Approx.}, 11(4):1--19, 2018.

\bibitem[BCM]{BCM}
T.~Bayraktar, D.~Coman, and G.~Marinescu.
\newblock Universality results for zeros of random holomorphic sections.
\newblock {\em Trans. Amer. Math. Soc.}, 373(6): 3765-3791, 2020.

\bibitem[B03]{Bernt03}
B.~Berndtsson,
\newblock Bergman kernels related to Hermitian line bundles over compact
complex manifolds.
\newblock Explorations in complex and Riemannian geometry, {\em Contemp. Math.}, 332, 1-17, 2003.
 
\bibitem[Ber09]{Berman}
R.~J. Berman.
\newblock Bergman kernels and equilibrium measures for line bundles over
  projective manifolds.
\newblock {\em Amer. J. Math.}, 131(5):1485--1524, 2009.

\bibitem[BBWN11]{BBWN}
R.~Berman, S.~Boucksom, and D.~Witt Nystr\"{o}m. 
\newblock Fekete points and convergence towards
equilibrium measures on complex manifolds. 
\newblock{\em Acta Math.}, 207: 1--27, 2011.

\bibitem[Bos18]{Bos}
L.~Bos.
\newblock  Fekete points as norming sets.
\newblock {\em Dolomites Res. Notes Approx.}, 11 Special Issue Norm Levenberg: 26--34, 2018.

\bibitem[C99]{Cat}
D.~Catlin.
\newblock The Bergman kernel and a theorem of Tian.
\newblock in Analysis and Geometry
in Several Complex Variables (Katata, 1997), {\em Trends Math.}, 1--23. Birkh\"auser,
Boston, MA, 1999.

\bibitem[C91]{Ch1} M.~Christ.
\newblock On the $\overline{\partial}$ equation in weighted 
$L^2$-norms in $\Bbb{C} ^1$
\newblock {\em J. Geom. Anal.} 3:193--230, 1991.

\bibitem[C13]{Ch2} M.~Christ.
\newblock Upper bounds for Bergman kernels associated to positive
line bundles with smooth Hermitian metrics.  
\newblock arXiv:1308.0062.

\bibitem[DLM06]{DLM} X.~Dai, K.~Liu, and X.~Ma, 
\newblock On the asymptotic expansion of Bergman kernel.
\newblock {\em J.\ Differential Geom.}, 72: 1--41, 2006.

\bibitem[FZ04]{FZ}
R.~Feng and S.~Zelditch.
\newblock
Median and mean of the supremum of $L^2$ normalized random.
\newblock{\em J. Funct. Anal.}, 266(8): 5085--5107, 2014. 

\bibitem[GH]{GH}
P.~Griffiths and J.~Harris
\newblock Principles of Algebraic Geometry, Wiley Classics Library, John
Wiley \& Sons, Inc., New York, 1994, Reprint of the 1978 original.

\bibitem[Le01]{Led}
M.~Ledoux.
\newblock The concentration of measure phenomenon 
\newblock {\em American Mathematical Soc.} Math. Surveys Monogr. Vol. 89, 2001.

\bibitem[LOC16]{LOC} N.~Lev and J.~Ortega-Cerd\`a
\newblock Equidistribution estimates for Fekete points on complex manifolds
\newblock {\em J. Eur. Math. Soc. (JEMS)}, 18(2): 425--464, 2016.

\bibitem[L12]{Li}
M.~Lifshits.
\newblock Lectures on Gaussian processes. In Lectures on Gaussian Processes.
\newblock {\em Springer, Berlin, Heidelberg}. 2012.

\bibitem[MM1]{MM07} X.\ Ma and G.\ Marinescu, 
{\em Holomorphic Morse Inequalities and Bergman Kernels}, 
Progress in Math., vol.\ 254, Birkh\"auser, Basel, 2007, xiii, 422 pp.

\bibitem[MM2]{MM15}
X.\ Ma and G.\ Marinescu, 
{\em Exponential estimate for the asymptotics of Bergman kernels},
Math.\ Ann.\ \textbf{362} (2015), no.\ 3-4, 1327--1347.

\bibitem[MOC10]{MOC}
J.~Marzo, and J.~Ortega-Cerd\'a.
\newblock Equidistribution of Fekete points on the sphere. 
\newblock {\em Constr. Approx.}, 32: 513--521, 2010.

\bibitem[T90]{Tia}
G.~Tian.
\newblock On a set of polarized K\"ahler metrics on algebraic manifolds.
\newblock {\em J. Differential Geom.}, 32: 99-130, 1990.

\bibitem[SZ99]{SZ}
B.~Shiffman and S.~Zelditch.
\newblock Distribution of zeros of random and quantum chaotic sections of
  positive line bundles.
\newblock {\em Comm. Math. Phys.}, 200(3):661--683, 1999.

\bibitem[SZ03]{SZ1}
B.~Shiffman and S.~Zelditch.
\newblock
Random polynomials of high degree and Levy concentration of measure
\newblock{\em Asian J. Math.}, 7(4): 627--646, 2003. 

\bibitem[SZ08]{SZ2}
B.~Shiffman and S.~Zelditch.
\newblock Number variance of random zeros on complex
manifolds.
\newblock {\em Geom. Funct. Anal.}, 18(4):1422--1475, 2008. 


\bibitem[S84]{Sun}
B.~S\"{u}ndermann.
\newblock
On projection constants of polynomial spaces on the unit ball in several variables
\newblock{\em Math. Z.}, 188(1): 111--117, 1984. 

\bibitem[Z98]{Zel}
S.~Zelditch.
\newblock Szeg\"{o} kernels and a theorem of Tian.
\newblock {\em Int. Math. Res. Notices} 331: 317--331, 1998.

\end{thebibliography}

\documentclass[12pt,leqno,oneside]{amsart}
\usepackage{mathrsfs,dsfont}
\usepackage{amsmath,amstext,amsthm,amssymb,bbm,mathtools,comment}
\usepackage{charter}
\usepackage{typearea}
\usepackage{hyperref}
\usepackage{color}

\usepackage{float}
\usepackage{tikz}
\usetikzlibrary{patterns}
\usepackage{graphicx}
\usepackage{hyperref}

\usepackage[width=6.4in,height=8.5in]
{geometry}

\def\bt{\begin{thm}}
\def\et{\end{thm}}
\def\bl{\begin{lem}}
\def\el{\end{lem}}
\def\bd{\begin{defn}}
\def\ed{\end{defn}}
\def\bc{\begin{cor}}
\def\ec{\end{cor}}
\def\bp{\begin{proof}}
\def\ep{\end{proof}}
\def\br{\begin{rem}}
\def\er{\end{rem}}

\newtheorem{thm}{Theorem}[section]
\newtheorem{prop}[thm]{Proposition}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{defn}[thm]{Definition}

\newtheorem{rem}[thm]{Remark}
\newtheorem{cor}[thm]{Corollary}

\numberwithin{equation}{section}

\title{Sampling Property of Fekete Points for \\ Random Holomorphic Sections}

\author{Turgay Bayraktar} 
\thanks{T.\ Bayraktar is supported by T\"{U}B\.{I}TAK grant ARDEB-3501/118F049 and Science Academy, Turkey-BAGEP}
\address{Faculty of Engineering and Natural Sciences, Sabanc{\i} University, \.{I}stanbul, Turkey}
\email{tbayraktar@sabanciuniv.edu}

\keywords{Fekete points, Sampling, Random holomorphic sections}
\subjclass[2000]{32U15, 32L05, 60D05}

\date{\today}

\begin{document}

\begin{abstract}
In this note, we obtain the growth order of Lebesgue constants for Fekete points associated with tensor powers of a positive line bundle. Moreover, by endowing the space of global holomorphic sections with a natural Gaussian probability measure we prove that Fekete points are sampling for random holomorphic sections.
\end{abstract}

\maketitle
\section{Introduction}
Let $X$ be a projective manifold of complex dimension $n$ and $L\to X$ be a holomorphic line bundle endowed with a positive Hermitian metric $h:=e^{-\varphi}$. We also denote the set of global holomorphic sections of $L$ by $H^0(X,L)$. For a fixed basis $\{S_1,\dots, S_N\}$ of $H^0(X,L)$ and $N$-tuple of points $x_1,\dots, x_N$ on $X$ where $N:=dim(H^0(X,L))$, we denote Vandermonde-type determinant by
$$\text{vdm}(x_1,\dots,x_N):=\det(S_i(x_j))_{1\leq i,j\leq N}$$
which defines a section of the pull-back line bundle $L^{\boxtimes N}$ over the manifold $X^N$. A configuration of points $(x_1,\dots, x_N)$ is called a \textit{Fekete configuration} if it maximizes the point-wise norm $|\det(S_i(x_j))|$ with respect to the Hermitian metric on $L^{\boxtimes N}$ induced by $h$. We remark that by compactness Fekete configurations always exist but they need not to be unique. 

Note that Fekete configurations form an interpolating array. Indeed, let's denote the \textit{Lagrange sections} (see \S \ref{lagrange} for definition) $\ell^k_j\in H^0(X,L^k)$ associated with a Fekete configuration $\mathcal{F}_k$ of order $k$ then one can define a projection
$$\pi_k:C(X,L^k)\to H^0(X,L^k)$$
$$\pi_k(s)(x)=\sum_{j=1}^{N_k}c_j\ell^k_j(x)$$ where $c_j:=\langle s(x_j), \ell_j(x_j)\rangle_{h^{\otimes k}}$ for $j=1,\dots,N_k$ and $C(X,L^k)$ denotes the set of continuous sections of the tensor power $L^k:=L^{\otimes k}$. We may endow both $C(X,L^k)$ and $ H^0(X,L^k)$ with the sup norm. Then the operator norm of the projection $\pi_k$ is given by 
$$\|\pi_k\|:=\Lambda_k=\max_{x\in X}\sum_{j=1}^{N_k}|\ell_j^k(x)|_{h^{\otimes k}}$$ which is called the \textit{Lebesgue constant} of order $k$. In particular, for every $s\in H^0(X,L^k)$ we have 
$$\|s\|_{\infty}:=\max_{x\in X}|s(x)|_{h^{\otimes k}}\leq \Lambda_k \max_{x_j\in \mathcal{F}_k}|s(x_j)|_{h^{\otimes k}}.$$ It follows from the definition of Lagrange sections that $\Lambda_k\leq N_k:=\dim(H^0(X,L^k))$. We also remark that the operator norm $\|\pi_k\|=\Lambda_k$ gives a bound on the distance (in the sup norm) between the interpolant and its best uniform approximation in $H^0(X,L^k)$. Indeed, it is easy to see that for $s\in C(X,L^k)$
$$\|s-\pi_k(s)\|_{\infty}\leq (1+\|\pi_k\|)\inf_{q\in H^0(X,L^k)}\|s-q\|_{\infty}.$$ 

Interpolating and sampling properties of Fekete configurations have been studied by many authors in various geometric settings (see \cite{Bernt03,BBWN,LOC}) and in the context of orthogonal polynomials (\cite{MOC,AOC, Bos} and references therein). Recall that in the context of polynomials on $\Bbb{C}^n$, Fekete configurations are defined to be those set of points that maximize the (absolute value of the) Vandermonde determinant on the given compact set $K\subset \Bbb{C}^n$. Fekete arrays provides often good (sometimes excellent) interpolation points. However, there are only a few cases $K$ for which precise location of Fekete points is known. For instance, Fej\'er showed that Fekete points of $K=[-1,1]$ are the end points and critical points of the $k^{th}$ Legendre polynomial. In this case, it is well-known that Lebesgue constants are of order $O(\log k)$ which is best possible (see eg.\ \cite{Sun}). Moreover, S\"{u}dermannn \cite{Sun} also showed that in the case of real sphere $K=S^n\subset \Bbb{R}^{n+1}$ Lebesgue constants are growing at order faster than $k^{\frac{n-1}{2}}$ for $n\geq 2$. Our first result gives asymptotic growth of Lebesgue constants in the current geometric setting:

\begin{thm}\label{th1}
Let $(X,\omega)$ be a compact K\"ahler manifold of dimension $n$ and $L\to X$ be a holomorphic line bundle endowed with a positive Hermitian metric $h$. The Lebesgue constant of order $k$ satisfy
\begin{equation}\label{lc}
\lim_{k\to \infty}\frac{\Lambda_k}{\dim H^0(X,L^k)}=1. 
\end{equation}
In particular, $\Lambda_k=k^n+O(k^{n-1})$ as $k\to\infty$.
\end{thm}

We say that an array $\Gamma_k:=\{x_j\}_{j=1}^{N_k}$ is \textit{separated} if there is a constant $\delta>0$ such that $d(x_i,x_j)\geq \frac{\delta}{\sqrt{k}}$ for $1\leq i\not=j\leq N_k$. We say that a separated array $\Gamma_k$ is \textit{sampling} for the pair $(L,h)$ if there is a uniform constant $C>0$ and $k_0\in\Bbb{N}$ such that

\begin{equation}\label{sampling}
\|s\|_{\infty}\leq C\max_{x_j\in\Gamma_k}|s(x_j)|_{h^{\otimes k}}
\end{equation}
 for every $s\in H^0(X,L^k)$ and $k\geq k_0$. 

In what follows we endow $H^0(X,L^k)$ with a natural Gaussian probability measure $Prob_k$ associated with the geometric data $(X,\omega,L,h)$ (see \S \ref{random} for details). One can also consider the product probability space $\prod_{k=1}^{\infty}(H^0(X,L^k), Prob_k).$ It follows from Theorem \ref{th1} that Fekete points $\mathcal{F}_k$ do not form a sampling array in the sense of (\ref{sampling}). However, they do form a sampling array for random holomorphic sections:
\begin{thm}\label{th2}
Let $(X,\omega)$ be a compact K\"ahler manifold and $L\to X$ be a holomorphic line bundle endowed with a positive Hermitian metric $h$. Assume that $H^0(X,L^k)$ is endowed with the Gaussian probability measure $Prob_k$ induced by the given geometric data. Then for each $\epsilon>0$ there exists $k_0\in \Bbb{N}$ and a uniform constant $C>0$ such that for $k\geq k_0$
\begin{itemize}
\item[(1)] there exists $ E_k\subset H^0(X,L^k)$ satisfying $Prob_k(E_k)=O(k^{-\epsilon})$

\item[(2)] $\|s\|_{\infty}\leq C \max_{x_j\in \mathcal{F}_k}|s(x_j)|_{h^{\otimes k}}$ for every $s\in H^0(X,L^k)\setminus E_k$.
\end{itemize} 
In particular, taking $\epsilon>1$, for almost every sequence of random holomorphic sections $\{s_k\}\in \prod_{k=1}^{\infty}(H^0(X,L^k), Prob_k)$ we have 
  $$\|s_k\|_{\infty}\leq C \max_{x_j\in \mathcal{F}_k}|s_k(x_j)|_{h^{\otimes k}}.$$
\end{thm}
Lev and Ortega-Cerd\'a \cite{LOC} recently studied sampling and interpolating properties of Fekete arrays. They proved that Fekete arrays $\mathcal{F}_k$ are separated i.e. $d(x_i,x_j)\gtrsim \frac{1}{\sqrt{k}}$ but Fekete points do not form a $L^2$-sampling array for the pair $(L,h)$. This was observed in \cite[Theorem 3]{LOC} by obtaining necessary density conditions for being $L^2$-sampling (respectively interpolating arrays) and showing that Fekete arrays have the critical density (see \cite[\S 7.2]{LOC}). In particular, our Theorem \ref{th1} gives precision on this by estimating the growth order of the the Lebesgue constants. Lev and Ortega-Cerd\'a also observed that there is no separated array $\Gamma_k$ that is simultaneously sampling and interpolating for the pair $(L,h)$ (see \cite[Theorem 5]{LOC}). Unlike in the deterministic case, Theorem \ref{th2} shows that Fekete arrays are sampling and interpolating for random holomorphic sections.

The outline of the paper is as follows: we review basic tools in K\"ahler geometry and Bergman kernel asymptotics in Section \ref{prelim}. We prove Theorem \ref{th1} in Section \ref{samplingsec}. We consider random holomorphic sections in Section \ref{random} and prove Theorem \ref{th2} therein. Finally, we obtain $L^2$-version of our results.

\section{Preliminaries}\label{prelim}
Let $(X,\omega)$ be a compact K\"ahler manifold of complex dimension $n$ and $L\to X$ be a holomorphic line bundle endowed with a smooth Hermitian metric $h=e^{-\varphi}$ where $\varphi=\{\varphi_{\alpha}\}$ is a local weight of the metric. The latter means that if $e_{\alpha}$ is a local holomorphic frame (i.e. a non-vanishing holomorphic section) for $L$ over an open set $U_{\alpha}$ then $|e_{\alpha}|_h=e^{-\varphi_\alpha}$ where $\varphi_{\alpha}\in \mathscr{C}^{\infty}(U_{\alpha})$ such that $\varphi_{\alpha}=\varphi_{\beta}+\log|g_{\alpha\beta}|$ and $g_{\alpha\beta}:=e_{\beta}/e_{\alpha}\in\mathcal{O}^*(U_{\alpha}\cap U_{\beta})$ are the transition functions for $L$. In what follows, we assume that the metric $h$ is positively curved this means the its curvature form ${\omega_h}_{|_{U_{\alpha}}}=i\partial \overline{\partial}\varphi_{\alpha}$ is a (globally well-defined) positive closed (hence, K\"ahler) form. We remark that under positivity assumption Kodaira's embedding theorem \cite{GH} implies that $X$ is a projective manifold. The K\"ahler form $\omega$ induces a distance function $d(x,y)$ which will be used to define the balls $B(x,r):=\{y\in X: d(x,y)<r\}$. 

The Hermitian metric $h$ induces a Hermitian metric $h^{\otimes k}$ on the $k^{th}$ tensor power $L^k:=L\otimes\dots\otimes L$ given by $|e_{\alpha}^{\otimes k}|_{h^{\otimes k}}:=|e_{\alpha}|_{h}^n$. Recall that a \textit{global holomorphic section} $s_k\in H^0(X,L^k)$ can be locally represented as $s_k=f_{\alpha}e_{\alpha}$ for some $f_{\alpha}\in\mathcal{O}(U_{\alpha})$ satisfying the compatibility condition $f_{\alpha}=f_{\beta}g_{\alpha\beta}^n$ on $U_{\alpha\beta}:=U_{\alpha}\cap U_{\beta}$. In what follows, for simplicity of notation we write $s_k=f_ke_k$ where $e_k:=e^{\otimes k}$ is a local non-vanishing holomorphic section defined as above. We remark that $H^0(X,L^k)$ is a finite dimensional complex vector space with 
\begin{equation}\label{dim}
\dim H^0(X,L^k)=k^n+O(k^{n-1}).
\end{equation} We refer the reader to the text \cite{GH} for detailed account of complex geometry tools used in this note. 

The geometric data given above allow us to define a scalar inner product on the vector space of global holomorphic sections $H^0(X,L^{\otimes k})$ via
\begin{equation}\label{inp}\langle s_1,s_2\rangle:=\int_X \langle s_1(x),s_2(x)\rangle_{h^{\otimes k}} dV
\end{equation} where $dV=\frac{\omega^n}{n!}$ is the fixed volume form on $X$. We also denote the induced norm by $\|s\|_2$. 
  
\textbf{Notation:} In the sequel, $\lesssim$ (resp. $\gtrsim$) indicates that the corresponding inequality holds up to a positive constant. 

The following construction of peak-sections is a well-known result goes back to Tian \cite{Tia}: 
\begin{lem}[H\"{o}rmander Peak Sections]\label{HorPeak}
Given a point $x\in X$ there exist $k_0\in\Bbb{N}$ such that for every $k\geq k_0$ there is a section $\phi_x\in H^0(X,L^k)$ such that
\begin{itemize}
\item[(i)] $|\phi_x(x)|_{h^{\otimes k}}=1$
\item[(ii)] $\|\phi_x\|_2^2=O(N^{-1}_k).$
\end{itemize}
\end{lem}

\subsection{Bergman Kernel Asymptotics}
Let $\{S_j^k\}$ be a fixed orthonormal basis (ONB) for $H^0(X,L^k)$ with respect to the inner product (\ref{inp}). Recall that \textit{Bergman kernel} for the Hilbert space $H^0(X,L^k)$ is the integral kernel of the orthogonal projection from $L^2$-space of global sections with values in $L^k$ onto $H^0(X,L^k).$ That is Bergman kernel satisfy the reproducing property
$$s(x)=\int_X\langle s(y),K_k(x,y)\rangle_{h^{\otimes k}} dV(y)$$ for $s\in H^0(X,L^k)$. Thus, it can be represented as a smooth section 
$$K_k(x,y)=\sum_{j=1}^{N_k}S^k_j(x)\otimes \overline{S_j^k(y)}$$
of the line bundle $L^{\otimes k}\boxtimes(L^{\otimes k})^*$ over $X\times X.$ Note that this representation is independent of the choice of ONB $\{S_j^k\}$. The point-wise norm of restriction of $K_k(x,y)$ to the diagonal is given by 
$$|K_k(x,x)|=\sum_{j=1}^{N_k}|S^n_j(x)|^2_{h^{\otimes k}}.$$ 
We remark that $|K_k(x,x)|$ is the dimensional density:
$$\int_X|K_k(x,x)|dV=\dim H^0(X,L^k)=N_k$$ for $k\in\Bbb{N}$. In the sequel, we will need the following standard result which follows from reproducing property of the Bergman kernel:

\begin{lem}\label{peak2}
Given a point $y\in X$ and $k\in\Bbb{N}$ there is a section $\phi_y\in H^0(X,L^k)$ such that
$$|\phi_y(x)|_{h^{\otimes k}}=|K_k(x,y)|\ \text{for}\ x\in X.$$
\end{lem}

Recall that celebrated Catlin-Tian-Zelditch \cite{Cat,Tia,Zel} theorem asserts that the Bergman kernel $K_k(x,y)$ has the following diagonal asymptotics:
\begin{thm}\label{neardiag}
Let $(X,\omega)$ be a compact K\"ahler manifold of complex dimension $n$ and $L\to X$ be a holomorphic line bundle endowed with a positive Hermitian metric $h$. Then 
\begin{equation} |K_k(x,x)|=\frac{(c_1(L,h))^n_x}{\omega_x^n}k^n+O(k^{n-1})\ \text{as}\ k\to\infty.
\end{equation}
\end{thm}
Off-diagonal asymptotics of the Bergman kernel has been considered by various
authors in different settings (see eg.\ \cite{Ch1,MM07,DLM, SZ2, BCM} and references therein). The following version is adapted from \cite{MM15}:
\begin{thm}\label{offdiag}
Let $(X,\omega)$ be a compact K\"ahler manifold of complex dimension $n$ and $L\to X$ be a holomorphic line bundle endowed with a positive Hermitian metric $h$. Then there exist $c>0$ and $k_0\in\Bbb{N}$ such that for any $r\in\Bbb{N}$ there exists $C_r$ such that for any $k\geq k_0$ and $x,y\in X$ we have
\begin{equation}
|K_k(x,y)|_{\mathcal{C}^r}\leq C_rk^{n+\frac{r}{2}}\exp(-c\sqrt{k}d(x,y))
\end{equation} 
\end{thm}
Here, the pointwise $\mathcal{C}^r$ seminorm $|K_k(x,y)|_{\mathcal{C}^r}$ of the section $K_k\in\mathcal{C}^{\infty}(X\times X, L^{\otimes k}\boxtimes (L^{\otimes k})^*)$ at the point $(x,y)\in X\times X$ is the sum of norms induced by $h$. 
\section{Sampling and Interpolating Properties of Fekete Points}\label{samplingsec}
\subsection{Fekete Configurations} Recall that a \textit{Fekete configuration} is a finite collection $\Gamma$ of points in $X$ maximizing the determinant in the interpolation problem. More precisely, let $N:=\dim H^0(X,L)$ and $\Gamma=(x_1,\dots, x_N)$ we consider the evaluation map
$$ev_{\Gamma}:H^0(X,L)\to \bigoplus_{j=1}^NL_{x_j}.$$ Then a Fekete configuration maximizes point-wise norm of the determinant of the evaluation map with respect to a (equivalently every) basis $\{S_1,\dots,S_N\}$ of $H^0(X,L)$. This means that writing $S_i=f_{ij}e_j$ for some local frame $e_j$ near the point $x_j$ a Fekete array maximizes the quantity 
$$|\det(S_i(x_j))|_h:=|\det(f_i(x_j))|e^{-\varphi_1(x_1)}\cdots e^{-\varphi_N(x_N)}.$$

\subsection{Lagrange Sections}\label{lagrange} Given a configuration $\Gamma=\{x_1,\dots,x_N\}$ of distinct points we define associated \textit{Lagrange sections} $\ell_j\in H^0(X,L\otimes L_{x_j}^*)$ for $j=1,\dots,N$ by
\begin{equation}
\ell_j(x):=\text{vdm}(x_1,\dots,x_{j-1},x,x_{j+1}\dots,x_N)\otimes \text{vdm}(x_1,\dots,x_j,\dots,x_N)^{-1}
\end{equation}
where $\text{vdm}(x_1,\dots,x_N):=\det(S_i(x_j))_{1\leq i,j\leq N}$. More precisely, writing $S_i=f_{ij}e_j$ for some local frame $e_j$ near the point $x_j$ we let $M=\big(f_i(x_j)e^{-\varphi_j(x_j)}\big)_{1\leq i,j\leq N}$ then we have
\begin{equation}
\ell_j(x)=\frac{1}{|\det M|}\sum_{i=1}^N(-1)^{i+j}M_{ij}S_i(x)
\end{equation}
where $M_{ij}$ denotes the determinant of the the submatrix of $M$ obtained by deleting $i^{th}$ row and $j^{th}$ column of $M$.
Note that 
$$|\ell_j(x_i)|_{h}=\delta_{ij}\ \text{for}\ 1\leq i,j,\leq N.$$ Thus, $\ell_j$ form a basis for $H^0(X,L)$ and for each $s\in H^0(X,L)$ we have $$s(x)=\sum_{j=1}^Ns(x_j)\otimes \ell_j(x)$$ 
We remark that if $\Gamma=\mathcal{F}$ is a Fekete array then then Lagrange sections have the property that $$\sup_{x\in X}|\ell_j(x)|_{h}\leq 1.$$
The following result was obtained in \cite[Lemma 3]{LOC} (cf. \cite[Lemma 3.1]{AOC}): 
\begin{prop}\label{sepFek}
Let $\mathcal{F}_k$ denote Fekete array of order $k$. Then there exists $k_0\in\Bbb{N}$ and $\delta>0$ such that
$$d(x_i^k,x_j^k)\geq \frac{\delta}{\sqrt{k}} $$ for distinct points $x_i^k,x_j^k\in\mathcal{F}_k$ and $k\geq k_0$.
\end{prop}

\subsection{Quantitative Equidistribution of Fekete Points}
Asymptotic distribution of Fekete points on complex manifolds is studied by Berman, Boucksom and Witt Nystr\"{o}m \cite{BBWN}. They proved that Fekete points are equidistributed with respect to (normalized) equilibrium measure induced by the geometric data $(X,L,h)$. Their result applies in setting of big line bundles equipped with a smooth metric. When $(L,h)$ is a positive line bundle Lev and Ortega-Cerd\'a \cite[Theorem 1]{LOC} obtained a quantitative version of this result by estimating the discrepancy between Fekete points and the limiting measure.  
  \begin{thm}\label{LOCthm}
Let $(X,\omega)$ be a compact K\"ahler manifold and $L\to X$ be a holomorphic line bundle endowed with a positive Hermitian metric $h$. Let $r_k$ be a sequence of positive real numbers satisfying $r_k\to\infty$ and $\frac{r_k}{\sqrt{k}}\to0$ as $k\to \infty$. Then
\begin{equation}
\frac{\#(\mathcal{F}_k\cap B(x,\frac{r_k}{\sqrt{k}}))}{\#\mathcal{F}_k}=(1+O(\frac{1}{r_k}))\frac{\int_{B(x,\frac{r_k}{\sqrt{k}})}(i\partial \overline{\partial}\varphi)^n}{\int_X(i\partial \overline{\partial}\varphi)^n}
\end{equation} 
uniformly $x\in X$.
\end{thm}
We remark that the statement of \cite[Theorem 1]{LOC} involves a fixed radii $\frac{r}{\sqrt{k}}$ for any $r>0$. One can prove Theorem \ref{LOCthm} by adapting approach of \cite{LOC} to the current setting . One of the key points is that the assumptions on $r_k$ ensures that 
$$\int_{X\setminus \Omega}|K_k(x,y)|dV(x)=o(\frac{1}{k^n})\ \text{as}\ k\to\infty$$  where $\Omega=B(x,\frac{r_k}{\sqrt{k}})$ (see eg.\ \cite{Bernt03, B4}). As the arguments are very similar to that of \cite[Theorem 1]{LOC} we omit the details.
\subsection{Lebesgue Constants}
Building upon the arguments in \cite{LOC} we prove the next result:
\begin{prop}\label{asamp}
For every $a>1$ the Fekete arrays $\mathcal{F}_{\lceil ak \rceil}$ form a sampling array for the pair $(L,h)$.
\end{prop}

\begin{proof}
We write $a=1+\epsilon$ for some $\epsilon>0$. Let $s\in H^0(X,L^k)$ then by compactness of $X$ we have $\|s\|_{\infty}=|s(x_0)|_{h^{\otimes k}}$ for some $x_0\in X$. By Lemma \ref{HorPeak} there exists $Q_{x_0}\in H^0(X,L^{\lceil \frac{\epsilon k}{2} \rceil})$ such that $|Q_{x_0}(x_0)|=1$ and $\|Q_{x_0}\|_2^2=O(N^{-1}_{\lceil \epsilon k \rceil})$. We set $$\phi:=s\otimes Q_{x_0}^2\in H^0(X,L^{\lceil(1+\epsilon)k\rceil}).$$
Note that $|\phi(x_0)|_{h^{\otimes k}}=|s(x_0)|_{h^{\otimes k}}=\|s\|_{\infty}$. 
Using Lagrange sections of level $\lceil(1+\epsilon)k\rceil$ we write  $$\phi(x_0)=\sum_{j= 1}^{N_{\lceil(1+\epsilon)k\rceil}}c_j \ell_j(x_0)$$ where $c_j:=\langle \phi(x_0),\ell_j(x_0)\rangle_{h^{\otimes \lceil(1+\epsilon)k\rceil}}$ we see that
\begin{eqnarray*}
\|s\|_{\infty} &=& |\phi(x_0)|_{h^{\otimes k}}\\
&=& |\sum_{j\geq 1}c_j \ell_j(x_0)|\\
&\leq& \sum_{x_j\in \mathcal{F}_{\lceil(1+\epsilon)k\rceil}} |s(x_j)|_{h^{\otimes k}} |Q_{x_0}(x_j)|^2\\
&\leq& \sup_{x_j\in \mathcal{F}_{\lceil ak \rceil}} |s(x_j)|_{h^{\otimes k}} \sum_{j=1}|Q(x_0))|^2\\
&\leq& C \sup_{x_j\in \mathcal{F}_{\lceil ak \rceil}}|s(x_j)|_{h^{\otimes k}}N_{\lceil(1+\epsilon)k\rceil}\int_X|Q_{x_0}|^2dV\\
&\leq& C_1 \sup_{x_j\in \mathcal{F}_{\lceil ak \rceil}}|s(x_j)|_{h^{\otimes k}}\frac{N_{\lceil(1+\epsilon)k\rceil}}{N_{\lceil\epsilon k\rceil}}
\end{eqnarray*}
where in the third inequality we used Plancherel-P\'olya type inequality \cite[Lemma 2]{LOC}. Finally, since $\frac{N_{\lceil(1+\epsilon)k\rceil}}{N_{\lceil\epsilon k\rceil}}$ is bounded by a uniform constant the assertion follows.
\end{proof}
Next, we prove Theorem \ref{th1}. Note that Theorem \ref{th1} implies that one can not improve Proposition \ref{asamp} to the case $a=1$.

\begin{proof}[Proof of Theorem \ref{th1}] We have already observed that $\Lambda_k\leq N_k:=\dim H^0(X,L^k)$. For fixed $0<\epsilon<1$ let $\rho_k:=\frac{\epsilon \log k}{(1-\epsilon)\sqrt{k}}$. For sufficiently large $k$, we will construct a section $s\in H^0(X,L^k)$ such that
\begin{equation}
(1-\epsilon)^nk^{\epsilon}N_k\lesssim \frac{\|s\|_{\infty}}{\max_{x_j\in\mathcal{F}_k}|s(x_j)|_{h^{\otimes k}}}
\end{equation} where the implied constant does not depend on $k$. To this end, fix $x\in X\setminus \mathcal{F}_k$. Then it follows from Theorem \ref{LOCthm} that there exists $k_0\in\Bbb{N}$ such that
$$\frac{\# \mathcal{F}_k\cap B(x,\rho_k)}{\#\mathcal{F}_k}= (1+O(\frac{1}{\log k}))\frac{\int_{B(x,\rho_k)}(i\partial \overline{\partial}\varphi)^n}{\int_X(i\partial \overline{\partial}\varphi)^n}\leq \epsilon^n$$ for $k\geq k_0$.

Next, we construct Lagrange sections $\tilde{\ell}_k\in H^0(X,L^{\lceil\epsilon k\rceil})$ associated with the set $\{x\}\cup\{x_j^k\in \mathcal{F}_k\cap B(x,\rho_k)\}$ satisfying $|\tilde{\ell}_k(y)|\leq1$ for $y\in X$ and 
$$|\tilde{\ell}_k(x)|=1\ \text{and}\ \tilde{\ell}_k(x_j^k)=0\ \text{for}\ x_j^k\in \mathcal{F}_k\cap B(x,\rho_k).$$

By Lemma \ref{peak2} there exists $\Phi_x\in H^0(X,L^{\lceil(1-\epsilon)k \rceil})$ such that $|\Phi_x(y)|=|K_{\lceil(1-\epsilon)k\rceil}(y,x)|$ for $y\in X$. Now, we let
$$s(y):=\tilde{\ell_k}(y)\otimes \frac{\Phi_x(y)}{|K_{\lceil(1-\epsilon)k\rceil}(x,x)|}\in H^0(X,L^k).$$
Note that $\|s\|_{\infty}= |s(x)|_{h^{\otimes k}}=1$. On the other hand, for $x_j^k\in  \mathcal{F}_k\cap B(x,\rho_k)$ we have $s(x_j^k)=0$. Moreover, for $x_j^k\in \mathcal{F}_k\setminus B(x,\rho_k)$ by Theorem \ref{neardiag} and Theorem \ref{offdiag} we have
$$|s(x_j^k)|_{h^{\otimes k}}\lesssim \frac{e^{-\sqrt{(1-\epsilon)k}\frac{\epsilon \log k}{(1-\epsilon)\sqrt{k}}}}{((1-\epsilon)k)^n}\lesssim \frac{1}{k^{\epsilon}(1-\epsilon)^nN_k}$$
where we used $N_k=\#\mathcal{F}_k=k^n(1+O(\frac1k))$ as $k\to\infty$. This implies that
$$(1-\epsilon)^nk^{\epsilon}N_k\lesssim \frac{\|s\|_{\infty}}{\max_{x_j^k\in\mathcal{F}_k}|s(f_j)|_{h^{\otimes k}}}.$$
Hence, the assertion follows by letting $\epsilon\to 0^+$.
\end{proof}

\section{Random Holomorphic Sections}\label{random} 
Let $\{S_j^k\}$ be a fixed orthonormal basis (ONB) for $H^0(X,L^k)$ with respect to the inner product (\ref{inp}). A \textit{Gaussian holomorphic section} is of the form
$$s(x):=\sum_{j=1}^{N_k}c_jS_j^k(x)$$ where $c_j$ are independent identically distributed (iid) complex Gaussian random variables of mean zero and variance $\frac{1}{N_k}$. This induces a $N_k$-fold product measure on $H^0(X,L^k)$ given by 
$$Prob_k:=(\frac{N_k}{\pi})^{N_k}e^{-N_k\|c\|^2}d\mathcal{L}(c)$$ where $d\mathcal{L}$ denotes the Lebesgue measure on $\Bbb{C}^{N_k}$ and $\|c\|$ denotes the Euclidean norm of the vector $\vec{c}=(c_j)_{j=1}^{N_k}$.
We remark that this Gaussian is characterized by the properties that the real random random variables $Re(c_j)$ and $Im(c_j)$ are independent with mean zero and satisfying $\Bbb{E}[c_j]=0$ together with
\begin{equation}\label{varnorm}
\Bbb{E}[c_ic_j]=0\ \  \&\ \ \Bbb{E}[c_i\overline{c_j}]=\frac{1}{N_k}\delta_{ij}
\end{equation} for all $i,j\in\{1,\dots,N_k\}.$ We also remark that this Gaussian ensemble is equivalent to the spherical ensemble induced by the norm $\|\cdot\|_2$. More precisely, denoting the unit sphere $SH^0(X,L^k):=\{s\in H^0(X,L^k):\|s\|_2=1\}$ we endow it with the Haar probability measure $\sigma_k$. Then using the spherical coordinates and integrating the radial variable out one sees that for $\mathcal{A}\subset H^0(X,L^k)$ 
$$Prob_k(\mathcal{A})=C_k\sigma_k(\Pi(\mathcal{A}))$$ where $\Pi:H^0(X,L^k)\setminus\{0\}\to SH^0(X,L^k)$ 
$$\Pi(s)(x)=\frac{s(x)}{\|s\|_2}$$ is the natural projection and $C_k>0$ is a dimensional constant.
 
Asymptotic distribution of zeros of random holomorphic sections was initiated by Shiffman and Zelditch \cite{SZ} (see \cite{B6} and references therein for non-Gaussian ensembles). They proved that (suitably normalized ) current of integrations along zeros of random holomorphic sections equidistributed with respect to the curvature form $\omega_h$ as $k\to \infty$. We refer the reader to the survey \cite{BCHM} and references therein for the current state of the art of this problem. More recently, Shiffman-Zelditch \cite{SZ1}  and Feng-Zelditch \cite{FZ} studied asymptotic growth of sup norms $\|s_k\|_{\infty}$ of random holomorphic sections. In particular, they proved the following result:

\begin{thm}[\cite{FZ}]\label{FZthm}
There exists uniform constants $C,c>0$ such that
$$Prob_k\{s\in H^0(X,L^k): \big|\|s\|_{\infty}-\sqrt{n\log k}\big|\geq \epsilon\}\leq Ck^{-c\epsilon^2}$$ for each $\epsilon>0$.
\end{thm}  

\subsection{Proof of Theorem \ref{th2}}
Now, we will prove Theorem \ref{th2}. For fixed $\epsilon>0$ it follows from Theorem \ref{FZthm} that there exists $C>0$ such that
\begin{equation}\label{sb}
Prob_k\{s\in H^0(S,L^k): \|s\|_{\infty}>C\sqrt{\log k}\}=O(k^{-c\epsilon^2}).
\end{equation}
Next, for each $x_j^k\in\mathcal{F}_k$ we define the Gaussian processes
$$Y_j^k:H^0(X,L^k)\to L^k_{x_j^k}$$
$$Y_j^k(s):=s(x^k_j)$$ for $j=1,\dots, \#\mathcal{F}_k$. 
First, we prove that for each $k$ we can find a subset $M_k\subset \mathcal{F}_k$ with positive density such that increment variance between $Y_j^k$'s for $x_j^k\in M_k$ are bounded from below:
\begin{lem}\label{lem1}
For each $0<\delta, c<1$ there exists $k_0\in\Bbb{N}$ such that for each $k\geq k_0$ there exist $M_k\subset \mathcal{F}_k$ satisfying $\#M_k\geq c\#\mathcal{F}_k$ and 
$$\Bbb{E}[|Y_i^k-Y_j^k|^2]\geq \delta$$ for $x_i^k,x_j^k\in M_k$ and $i\not=j$.
\end{lem}
\begin{proof}
Fix $0<\delta<1$. As before we write $S^k_j=f_j^ke_k$ for a local frame $e_k$ near the point $x_j^k$. Then by Theorem \ref{neardiag} and Theorem \ref{offdiag} we obtain
\begin{eqnarray}
\Bbb{E}[|Y^k_i-Y^k_j|^2] &=& \Bbb{E}|\big(\sum_{\ell=1}^{N_k}c_{\ell}^kf_{\ell}^k(x_i^k)\big)e^{-n\phi(x_i^k)}-\big(\sum_{\ell=1}^{N_k}c_{\ell}^kf_{\ell}^k(x_j^k)\big)e^{-n\phi(x_j^k)} \nonumber \\
&=&\frac{1}{N_k}\big(K_k(x_i^k,x_i^k)+K_k(x_j^k,x_j^k)-2Re(K_k(x_i^k,x_j^k))\big)\nonumber \\
&\gtrsim& 2+O(\frac1k)+O(k^n\exp(-\sqrt{k}d(x_i^k,x_j^k))) \label{eq3}. 
\end{eqnarray}
 Let $r_k>0$ with the property $r_k\to\infty$ and $\frac{r_k}{\sqrt{k}}\to0$ as $k\to \infty$. Then by (\ref{eq3}) $d(x_i^k,x_j^k)\geq r_k$ implies that
\begin{equation}
\Bbb{E}[|Y^k_i-Y^k_j|^2]\geq \delta
\end{equation}
for $k$ sufficiently large.

Next, we define $M_k$ to be the maximal set of points in $\mathcal{F}_k$ whose elements' pairwise distance greater than $\frac{r_k}{\sqrt{k}}$. We will show that for every $0<c<1$ we have $\#M_k\geq c\#\mathcal{F}_k$ for sufficiently large $k$. Indeed, for each $x_j^k\in \mathcal{F}_k$ by Theorem \ref{LOCthm} and $\mathcal{F}_k=N_k=k^n(1+O(\frac1k))$ we have
\begin{equation}\label{number}
\frac{\#\mathcal{F}_k\cap B(x_j^k,\frac{2r_k}{\sqrt{k}})}{\#\mathcal{F}_k}= (1+O(\frac{1}{r_k}))\frac{\int_{B(x_j^k,\frac{2r_k}{\sqrt{k}})}(i\partial \overline{\partial}\varphi)^n}{\int_X(i\partial \overline{\partial}\varphi)^n}= (1+O(\frac{1}{r_k}))O(\frac{2r_k}{\sqrt{k}})^{2n}.
\end{equation} 
Now, we let $C_k$ denote the least number of balls of the form $B(x_j^k,\frac{2r_k}{\sqrt{k}})$ with $x_j^k\in \mathcal{F}_k$ needed to cover $\mathcal{F}_k$. Note that  $\#M_k\geq C_k. $ On the other hand, by (\ref{number}) we have $C_k\gtrsim \frac{k^n}{r_k^{2n}}$. Thus, the assertion follows. 
\end{proof}

We may write each $Y_j^k=Z_{1,j}^k+iZ_{2,j}^k$ where $Z_{i,j}^k$ are independent real centered Gaussian processes. Observe that
\begin{equation}\label{eq4}
\sup_X|Y_j^k|\geq \sup_X\sqrt{(Z_{1,j}^k)^2+(Z_{2,j}^k)^2}\geq \frac{1}{\sqrt{2}}\sup_X|Z_{1,j}^k+Z_{2,j}^k|\geq \frac{1}{\sqrt{2}}\sup_X (Z_{1,j}^k+Z_{2,j}^k). 
\end{equation}
We also remark that by independence we have
\begin{equation}\label{eq5}
\Bbb{E}[(Z_{1,j}^k+Z_{2,j}^k-Z_{1,l}^k-Z_{2,l}^k)^2]=\Bbb{E}[|Y^k_i-Y^k_j|^2].
\end{equation}
Next, we will use Sudokov's minoration theorem (see \cite[Theorem 10.4]{Li}):

\begin{thm}\label{sudakov}
Let $X(t), Y(t)$ be two centered (real) Gaussian processes parametrized by a common set $T$. Assume that
$$\Bbb{E}[(Y(t)-Y(s))^2]\geq \Bbb{E}[(X(t)-X(s))^2]\ \ \forall s,t\in T.$$
Then $$\Bbb{E}[\sup_{t\in T}Y(t)]\geq \Bbb{E}[\sup_{t\in T}X(t)].$$
\end{thm} 
We will also use the following well-known lower bound for independent Gaussian processes (see. \cite[\S10.3]{Li}):
\begin{lem}\label{lem2}
Let $X_j$ be independent centered (real) Gaussian random variables for $j=1,\dots, N$. Assume that $$\min_{1\leq j\leq N}\Bbb{E}[X_j^2]\geq \sigma^2>0.$$ Then there exists $c>0$ such that
$$\Bbb{E}[\max_{1\leq j\leq N} X_j]\geq c\sigma\sqrt{\log N}.$$
\end{lem}

Thus, it follows from Theorem \ref{sudakov} and Lemma \ref{lem1} together with (\ref{eq4}) \& (\ref{eq5}) and Lemma \ref{lem2} that there exists $T>0$ such that
\begin{equation}
\Bbb{E}[\max_{x_j^k\in \mathcal{F}_k}|s(x_j^k)|_{h^{\otimes k}}]\geq T\sqrt{n\log k}+O(\frac1k)\ \text{as}\ k\to \infty.
\end{equation} 

On the other hand, fluctuations of $\Bbb{E}[\max_{x_j\in \Gamma_k}|s(x_j^k)|_{h^{\otimes k}}]$ is governed by the individual Gaussian process $|s(x_j^k)|_{h^{\otimes k}}$. This is a consequence of the following result in the context of Gaussian processes (see \cite[\S 7.1]{Led} ):

\begin{lem}[\cite{Led}]\label{Led}
Let $(X_t)_{t\in I}$ be centered (real) Gaussian processes with finite index set $I$. Assume that $\sigma^2:=\sup_{t\in I}(\Bbb{E}[X_t^2])<\infty$. Then for each $\epsilon>0$ 
$$Prob[\big|\sup_{t\in I}X_t-\Bbb{E}[\sup_{t\in I}X_t]\big|\geq \epsilon]\leq2\exp(-\frac{\epsilon^2}{2\sigma^2}).$$
\end{lem}
We remark that by Theorem \ref{neardiag} 
$$\Bbb{E}[|Y_j^k|^2|=\frac{1}{N_k}|K_k(x_j^k,x_j^k)|\lesssim\frac{(c_1(L,h))^n_{x_j^k}}{\omega_{x_j^k}^n}+O(\frac1k)=O(1).$$
Hence, applying Lemma \ref{Led} we obtain:
\begin{lem}\label{lem3}
For each $\epsilon>0$ there exists constants $A,b>0$ such that 
$$Prob_k\{s\in H^0(X,L^k): \big|\max_{x_j^k\in \mathcal{F}_k}|s(x_j^k)|_{h^{\otimes k}}-\Bbb{E}[\max_{x_j^k\in \mathcal{F}_k}s(x_j^k)|_{h^{\otimes k}}]\big|>\epsilon\sqrt{\log k}\} \leq Ak^{-b\epsilon^2}.$$
\end{lem}

Next, we define 
$$\mathcal{A}_k:=\{s\in H^0(S,L^k): \|s\|_{\infty}>C\sqrt{\log k}\}$$
 $$\mathcal{B}_k:=\{s\in H^0(X,L^k): \max_{x_j^k\in\Gamma_k}|s(x_j^k)|_{h^{\otimes k}}<(T+\epsilon)\sqrt{\log k}\}$$
and $E_k:=\mathcal{A}_k\cup\mathcal{B}_k$. Then by Lemma \ref{lem1} and Lemma \ref{lem3} we have
$$Prob_k(E_k)=O(k^{-\epsilon^2})$$ and
for $s\in H^0(X,L^k)\setminus E_k$ we have
$$\|s\|_{\infty}\leq \frac{C}{T+\epsilon}\max_{x_j^k\in \mathcal{F}_k}|s(x_j^k)|_{h^{\otimes k}}.$$

Finally, the last assertion in the Theorem follows from Borel-Cantelli lemma by putting $\epsilon>1$. This finishes the proof of Theorem \ref{th2}.

\subsection{$L^2$-Sampling for Random Holomorphic Sections} Next, we obtain a generalization of Theorem \ref{th2} for $L^2$-sampling arrays. Recall that a separated array $\Gamma_k$ is called an $L^2$-sampling for $(L,h)$ is there exists $k_0\in\Bbb{N}$ and uniform constants $A,B>0$ such that 
$$Ak^{-n}\sum_{\lambda\in \Gamma_k}|s(\lambda)|^2_{h^{\otimes k}}\leq \int_X|s(x)|^2_{h^{\otimes k}}dV\leq Bk^{-n}\sum_{\lambda\in \Gamma_k}|s(\lambda)|^2_{h^{\otimes k}}.$$

\begin{thm}\label{L2}
Let $(X,\omega)$ be a compact K\"ahler manifold and $L\to X$ be a holomorphic line bundle endowed with a positive Hermitian metric $h$. Assume that $H^0(X,L^k)$ is endowed with the Gaussian probability measure $Prob_k$ induced by the given geometric data. Then for each $\epsilon>0$ there exists $k_0\in \Bbb{N}$ and a uniform constants $A,B>0$ such that for $k\geq k_0$
\begin{itemize}
\item[(1)] there exists $ E_k\subset H^0(X,L^k)$ satisfying $Prob_k(E_k)=O(k^{-\epsilon})$

\item[(2)] $Ak^{-n}\sum_{\lambda\in \mathcal{F}_k}|s(\lambda)|^2_{h^{\otimes k}}\leq \int_X|s(x)|^2_{h^{\otimes k}}dV\leq Bk^{-n}\sum_{\lambda\in \mathcal{F}_k}|s(\lambda)|^2_{h^{\otimes k}}$ for every $s\in H^0(X,L^k)\setminus E_k$.
\end{itemize} 
In particular, taking $\epsilon>1$, for almost every sequence of random holomorphic sections $\{s_k\}\in \prod_{k=1}^{\infty}(H^0(X,L^k), Prob_k)$ we have 
  $$Ak^{-n}\sum_{\lambda\in \mathcal{F}_k}|s(\lambda)|^2_{h^{\otimes k}}\leq \int_X|s(x)|^2_{h^{\otimes k}}dV\leq Bk^{-n}\sum_{\lambda\in \mathcal{F}_k}|s(\lambda)|^2_{h^{\otimes k}}.$$
\end{thm}
Since the proof is very similar to that of Theorem \ref{th2} we provide only key ingredients and leave some details to the reader: 
\begin{proof}
Since $\mathcal{F}_k$ is separated the left inequality in (2) is a consequence of Plancherel-P\'olya type inequality \cite[Lemma 2]{LOC}. 

It follows from \cite[Eq. 3.12]{B8} that there exists $C_1c>0$ such that
\begin{equation}
Prob_k\{s\in H^0(X,L^k): \big|\int_X|s|^2_{h^{\otimes k}}dV-\int_Xi\partial\overline{\partial}\varphi\big|>\epsilon\}\leq C_1e^{-c\epsilon^2}
\end{equation} for sufficiently large $k$.

On the other hand, for each small $\epsilon>0$ we may find $M_k\subset \mathcal{F}_k$ as in proof of Theorem \ref{th2} such that $\# M_k\geq (1-\epsilon)\mathcal{F}_k$ and $d(x_i^k,x_j^k)\geq \frac{r_k}{\sqrt{k}}$ for $i\not= j$ where $r_k$ as in the proof of Theorem \ref{th2}. Then by Theorem \ref{offdiag} the covariances satisfy
\begin{equation}\label{varest}
\Delta_{ij}:=\Bbb{E}[Z^k_i\overline{Z}^k_j]=\frac{1}{N_k}|K_k(x_i^k,x_j^k)|\lesssim \exp(-c\sqrt{k}d(x_i^kx_j^k))\leq\exp(-cr_k) \ \text{for}\ i\not=j 
\end{equation}
where $Z^k_j(s):=s(x_j^k)$ and $x_j^k\in M_k$. We also remark that by (\ref{varnorm})
\begin{equation}\label{varest2}
\Delta_{jj}=\Bbb{E}[|Z_j^k|^2]=\Bbb{E}[|s(x_j^k)|^2_{h^{\otimes k}}]=\frac{1}{N_k}|K_k(x_j^k,x_j^k)|=O(1). \end{equation}
Note that $X_j^k(s):=|Z_j^k(s)|^2=|s(x_j^k)|^2_{h^{\otimes k}}$ is exponentially distributed random variables with $\Bbb{E}[X_j^k]=\frac{1}{N_k}|K_k(x_j^k,x_j^k)|$. Then it follows from (\ref{varest}) and (\ref{varest2}) and a Chernoff type bound for sums of weakly dependent random varibles we obtain
\begin{equation}
Prob_k\big\{s\in H^0(X,L^k): \big|\frac{1}{\# M_k}\sum_{x_j^k\in M_k}|s(x_j^k)|_{h^{\otimes k}}^2 - \frac{1}{(\# M_k) N_k}\sum_{x_j^k\in M_k}|K_k(x_j^k,x_j^k)|\big|<\epsilon\big\}\leq C_2e^{-c\epsilon^2}
\end{equation} for some $C_2>0$ endependent of $M_k$.
This in turn implies that there exists $E_k\subset H^0(X,L^k)$ such that $Prob(E_k)=O(k^{-\epsilon^2})$ and for $s\in H^0(X,L^k)\setminus E_k$ and sufficiently large $k$
\begin{equation}
\frac{\int_X|s|^2_{h^{\otimes k}}dV}{\frac{1}{\#M_k}\sum_{x_j^k\in M_k}|s(x_j^k)|^2_{h^{\otimes k}}}\leq C_3
\end{equation}
where $C_3>0$ independent of $M_k$. Thus, using $N_k=k^n+O(k^{n-1})$ and letting $\epsilon\to 0^+$ the assertion follows.
\end{proof}

\end{document}